OpenStudy (anonymous):

IS THIS EQUATION TRUE FALSE OR OPEN 4Y+8=6Y+3

OpenStudy (anonymous):

It should be false, but I'm not sure.

OpenStudy (anonymous):

could i help?

OpenStudy (anonymous):

it is open i am pretty sure

undeadknight26 (undeadknight26):

I believe this is true.

OpenStudy (anonymous):

@undeadknight26 you are wrong look at this http://en.wikipedia.org/wiki/Open_sentence

OpenStudy (anonymous):

it is not 2y+1=2y+1

undeadknight26 (undeadknight26):

If it was false there would be no answer...unless thats not true...

OpenStudy (anonymous):

4y+8=6y+3 therefore 6=2y

OpenStudy (solomonzelman):

My bad! Sorry!

OpenStudy (anonymous):

@SolomonZelman you cannot factor it out like that

OpenStudy (solomonzelman):

I know, I've been up since 5 am, sorry about that! Of course not, yoou're right it is open!!!!!!!!!!!!!

OpenStudy (anonymous):

it is an equation .. both sides have to be solved simultaneously @SolomonZelman

OpenStudy (anonymous):

hehe ok

OpenStudy (solomonzelman):

I am just tired... sorry again!